A $618,478 warehouse is being purchased by your company. The deal requires a down payment of 109,363 with the remainder of the purchase price paid over 20 years, payments in advance. The annual interest rate applicable is 6.58. What monthly payment in advance will be required to pay off this load? Key in your answer as a positive number.
Answer #1

Loan amount (PV) = 618,478 - 109,363 = 509,115

FV = 0

rate = 6.58/12

N = 20*12

use PMT funciton in Excel with type as 1

monthly payment = 3,799.01